OpenStudy (rocklionmba):

Is square root 1 minus sine squared theta = cos Θ true? If so, in which quadrants does angle Θ terminate?

OpenStudy (rocklionmba):

@One098

OpenStudy (anonymous):

remember that \(cos^2(\theta) + sin^2(\theta) = 1\), so \(cos^2(\theta)=1-sin^2(\theta)\)

OpenStudy (anonymous):

\(cos(\theta) = \sqrt{1-sin^2(\theta)}\)

OpenStudy (rocklionmba):

so then \[\cos(\theta) =\sqrt{\cos ^{2}(\theta)+\sin ^{2}(\theta)-\sin ^{2}(\theta)}\]

OpenStudy (freckles):

\[\sqrt{\cos^2(\theta)}=\cos(\theta) \text{ if } \cos(\theta)>0 \\ \sqrt{\cos^2(\theta)}=-\cos(\theta) \text{ if } \cos(\theta)<0\]

OpenStudy (anonymous):

yeah... but why you would write like this?

OpenStudy (anonymous):

yes @freckles is correct

OpenStudy (rocklionmba):

the two sins cancel out, so its just \[\cos (\theta)=\sqrt{\cos ^{2}(\theta)}\]

OpenStudy (rocklionmba):

I write like that to make it eaiser to see

OpenStudy (anonymous):

\(cos(\theta)\) is negative in the second and third quadrants

OpenStudy (rocklionmba):

I can tell its true because its squared and has a square root, and then assuming "terminate" means negative, which makes you right